OpenStudy (anonymous):

Factor x^3-5x^2-25x+125

OpenStudy (anonymous):

5x(x^2-x-5+25)

OpenStudy (welshfella):

- no - thats gives 5x^3

OpenStudy (welshfella):

x+ 5 or x-5 might be a factor find f(5) and see if it = 0

OpenStudy (anonymous):

So (x+5)(x-5)(x-5)

OpenStudy (anonymous):

What type of roots of those?

OpenStudy (michele_laino):

we can factor out x^2 between the first two terms, and we can factor out -25x between the third and the fourth term, so we get this: \[\Large {x^3} - 5{x^2} - 25x + 125 = {x^2}\left( {x - 5} \right) - 25x\left( {x - 5} \right)\]

OpenStudy (welshfella):

OOOH - did you guess that?? (x-5) is a factor because f(5) = 0

OpenStudy (anonymous):

No. I just did (x^2-25)(x-5)

OpenStudy (welshfella):

well you are correct

OpenStudy (anonymous):

What type of roots are those?

OpenStudy (anonymous):

And thanks youu

OpenStudy (michele_laino):

oops... we can factor out 25 between the third and the fourth terms, not 25 x: \[\Large {x^3} - 5{x^2} - 25x + 125 = {x^2}\left( {x - 5} \right) - 25\left( {x - 5} \right)\]

OpenStudy (welshfella):

right

OpenStudy (welshfella):

- you beat me too it!!

OpenStudy (anonymous):

So would the roots be -5,-5,5 or -5,5,5?

OpenStudy (welshfella):

the roots are 5 and -5 root 5 is a duplicate

OpenStudy (anonymous):

Ok but since the polynomial is 3 there has to be 3 roots?

OpenStudy (welshfella):

- correction - that was rubbish the roots are 5 , 5 and -5

OpenStudy (anonymous):

How?

OpenStudy (welshfella):

yes -5 and 5 duplicity 2

OpenStudy (welshfella):

factors are (x + 5)(x - 5)(x - 5) the fist gives x = -5 and the other 2 give x = 5

OpenStudy (anonymous):

Ohhh<3

OpenStudy (anonymous):

Thanks! :)

OpenStudy (welshfella):

no 3 roots - because there are 2 5's

OpenStudy (welshfella):

the graph will look a bit like